OpenStudy (anonymous):

what horizontal force will accelerate a 400kg crate at 1 m/s/s across a factory floor against a friction force half its weight

OpenStudy (anonymous):

Taking 'g' to be 10ms^-2, therefore weight = 400*10= 4000N Frictional force= 2000N Using netwon 2nd law 'R.F= ma', we have: X-2000= 400*1 X= 400+2000= 2400N